OpenStudy (anonymous):

multipy and simplify (4y-1)(8y+3)

mathslover (mathslover):

(4y-1)(8y+3) \[4y(8y)+4y(3)-1(8y)-1(3)\] \[32y^2+12y-8y-3\] \[32y^2+4y-3\]

mathslover (mathslover):

remember the distributive property (a-b)(c-d) ac-ad-bc+bd

OpenStudy (goformit100):

a trick to solve these type of questions.....remember it :- (a+b)(c+d) =(ac+ad)+(bc+bd) =ANSWER

OpenStudy (anonymous):

thank you ok!!